Present: Andre, Luigi, Heiko, Sebastian D., Alena, Alex, Annie, Benni, David, Florian, Dohun, Frank, Jan. Konstantin, Sebastian P., Thomas.

Minute taker: Luigi

Andre: new corona-rules in Germany should not change the situation in the department.

-Round table

Heiko: MP10 wafers sent to OPTIM, no concerns from them. Bochum confirms they have vacuum chuck but no UV-lamp. OPTIM will cure Bochum's wafer separately.

Sebastian D.: no news. Results shown today by Nick.

Alena: still working. Moved state machine from software to firmware. Seems to be working. Andre: who is looking at Run2020's small factor chips? Alena: don't know. So far she knows about hit counter chips.

Alex: worked with Luigi. Would like to start working on Run2020 next week. Andre: skype meeting on Monday.

Annie: worked on software to analyze data from testbeam from Vadym. Next week should be possible to start. Andre: interesting thing to tilt the planes of the telescope like Fermilab does. Needs more studies and works better with larger depletion region (larger resistivity).

Benni: start working again next week.

David: lot of data form testbeam still to analyze. Reading theses. Run2020 sensors were not optimized during testbeam. Expected efficiency low (>90% but <99%). In vacation from next week.

Florian: short scan of delay and amplitude for MP10. Probably asymmetrical. Tried new oscilloscope, maybe not enough resolution. David: these scopes have drivers compatible with our setup and faster. Jan: have you asked Stefan for DRS ports? Not possible, too slow for our signals.

Dohun: writing.

Frank: not much. Reported this morning. Worked on He system for the Swiss proposal. PSI is restricting to 50% of the people, beam still on.

Jan: feeling better at home. Applied for asylum in HD. Still figuring out. Haven't booked a flight back. At the moment can stay home only.

Konstantin: writing.

Sebastian P.: just handed in the thesis. Andre: can you keep working on testbeam data? No, busy semester.

Thomas: currently constructing the thermomecanical Mockup. Struggling with mechanics. Hopefully bringing it to Brugg tomorrow. 4/108 chips broken. 2 today because of design mistake, too tight for tolerances. Learning. On Monday will work on 6-chip PCB, Ralf is available.

Luigi: still debugging the probecard. Some progress but not definitive. Adapter card design reviewed with Dirk Gottschalk as regards the connections, will be reviewed with workshop for the mechanics. At the moment, it is 8.5 cm long.